# Adapting Vision-Language Models for Evaluating World Models

### v1

@@ 4 Experiments
Figure 2: Comparison of universe and baseline models on Action and Character Recognition. Left: universe outperforms all baselines on AR. Right: On CR, it ranks third, behind models with either full vision encoder tuning or task-specific training with greater supervision. Trained under a unified protocol with minimal parameter updates (0.07%) and reduced per-task data, universe delivers strong performance across both tasks, highlighting its efficiency and generalization.
Baselines. We compare universe against two classes of baselines: (i) Zero-shot VLMs:Seven off-the-shelf models, including VideoLLaMA3 (2B, 7B) [14] and three PaliGemma models: version 1 (3B) and version 2 (3B and 10B) [10, 91], evaluated without domain adaptation using an 8-frame visual context window. [^1] We also experimented with CLIPScore-based evaluation (Appendix F.2); results underperformed relative to selected baselines and were constrained to predefined candidate sets, further underscoring the need for model adaptation. (ii) Fine-tuned PaliGemma 2:Variants adapted via full, partial, and parameter-efficient tuning. This backbone is selected based on a sweep over PaliGemma variants, using zero-shot performance as a guide (Appendix F.1). The adaptation space includes 8 primary baselines: (i) Single-component fine-tuning: tuning only the vision encoder ($\mathcal{F}_{\text{V}}$), the multimodal projector ($\mathcal{F}_{\text{P}}$), or the language head ($\mathcal{F}_{\text{L}}$); (ii) Two-component fine-tuning: jointly tuning pairs of components—$\mathcal{F}_{\text{V+P}}$, $\mathcal{F}_{\text{V+L}}$, and $\mathcal{F}_{\text{P+L}}$; (iii) Full-model fine-tuning: tuning all components simultaneously ($\mathcal{F}_{\text{V+P+L}}$); (iv) LoRA-based tuning: Parameter-efficient adaptation with rank $r=8$, selected after observing minimal performance variation across $r\in\{8,16,32,48,64\}$ (see Appendix F.3 for details). All models are trained using 8-frame clips and a single epoch.
Results. Figure 2 (left, center) summarizes performance across Action Recognition (AR) and Character Recognition (CR). Zero-shot models perform poorly across both tasks: VideoLLaMA3 variants fall below 12.7% on AR and 6.4% on CR; PaliGemma variants reach up to 29.7% on AR and 17.2% on CR. These results confirm that general-purpose VLMs lack the temporal grounding and domain-specific semantics required for structured rollout evaluation. In contrast, universe achieves strong performance—outperforming all models on AR and ranking third on CR. The top two CR models either fine-tune the full vision encoder, requiring updates to $\sim$400M parameters, or fine-tune the multimodal projector with 5 $\times$ more CR supervision, each trained in isolation on a single task and prompt format. universe, by comparison, tunes only the 2.66M-parameter projector (0.07% of the model), under a unified protocol spanning both tasks, all prompt formats, and reduced per-task supervision. Its performance under these constraints underscores the efficiency and generality of our adaptation strategy for temporally grounded evaluation.
@@ 5 Analysis
[…]
@@
Figure 3: Exact Match accuracy over training epochs for Action Recognition (AR, top) and Character Recognition (CR, bottom). AR improves gradually with supervision, while CR converges to high accuracy with minimal training.
[…]
@@
[…]
@@
[…]
@@ 6 Evaluating World Model Rollouts with universe
We evaluate universe as an automated evaluator of world model rollouts using the WHAM benchmark [56], which provides pretrained world models and an evaluation set. Our analysis focuses on two axes: (i) in-domain accuracy, measured on Skygarden—the environment used during fine-tuning—and (ii) generalization to six unseen environments.
We compare universe’s predictions on samples from two world models: (i) WHAM-140M, trained on Skygarden with lower-quality rollouts ($128\times 128$ resolution), and (ii) WHAM-1.6B, trained on a diverse environment suite with higher-resolution output ($300\times 180$). We prepare 30 rollouts for each model-environment pair, yielding a total of 240 rollouts. Each rollout is segmented into 14-frame clips and paired with six natural language questions, following our evaluation protocol. universe answers each question using majority voting over five greedy decoding samples. Human annotators then rate each response on a four-point ordinal scale: Correct, Partially Correct, Incorrect, and Unclear. Each response is independently rated by two annotators; in cases of disagreement, a third annotator serves as adjudicator. Inter-annotator agreement is quantified using Cohen’s $\kappa$. Full details of the annotation protocol are provided in Appendix E.
Results. Figure 8 summarizes graded accuracy across models and environments. We observe a significant performance gap between rollouts from WHAM-140M and WHAM-1.6B. Despite being in-domain, WHAM-140M yields lower evaluation accuracy likley due to a resolution mismatch: its $128\times 128$ frames are upsampled to $224\times 224$ for universe input. In contrast, WHAM-1.6B produces higher-quality inputs, enabling more reliable evaluation. On WHAM-1.6B Skygarden rollouts, universe achieves 75.02% graded accuracy on AR and 90.00% on CR. When applied to six previously unseen environments, AR accuracy remains relatively stable with the lowest accuracy of 66.11% for environment G, and highest accuracy of 83.33% for environment E. Overall, the results indicate strong generalization, especially for identity-focused recognition tasks. Cohen’s $\kappa$ scores further reflect the interpretability of universe’s outputs: agreement is substantial overall ($\kappa=0.73$), with the lowest in Environment G ($\kappa=0.59$) and highest in Environment A on WHAM-1.6B rollouts ($\kappa=0.91$). These results highlight that universe remains aligned with human judgments.
Figure 8: Graded accuracy of universe across rollouts from WHAM-140M (Env. A) and WHAM-1.6B (Envs. A–G). Performance improves markedly with higher-fidelity rollouts from WHAM-1.6B, even in out-of-domain settings (B–G). Cohen’s $\kappa$ (above bars) reflects inter-rater agreement.
@@ 7 Conclusion
In this paper, we investigate the use of Vision-Language Models (VLMs) as automated evaluators for world model rollouts, addressing the fundamental challenge of fine-grained, temporally grounded evaluation. We introduce a structured evaluation protocol centered on action and character recognition tasks across binary, multiple-choice, and open-ended formats. To support this, we propose universe, a unified method for adapting VLMs to this setting through mixed supervision, efficient frame sampling, and lightweight fine-tuning. Our large-scale study demonstrates that universe matches the performance of task-specific baselines using a single checkpoint and aligns closely with human judgments, establishing it as a scalable, semantics-aware evaluator for evaluating world models, particularly when explicit ground truth is unavailable or costly to obtain.
Limitations While universe performs well in simulation, its generalization beyond simulated environments remains an open challenge. The evaluation protocol targets two fidelity axes, which, while comprehensive, omit higher-order reasoning over goals, causality, and multi-agent dynamics. Our experiments focus on short to medium context lengths; scaling to long-horizon rollouts remains an open challenge, especially under limited supervision. Although compute-efficient, training could be further improved with adaptive curricula or progressive tuning. Finally, like all pretrained VLMs, universe may reflect dataset biases and underperform on rare or ambiguous behaviors.
Supplementary Materials
[…]
@@ Appendix A Broader Impact
[…]
This capability has direct implications for high-impact domains such as neural game engines [56, 35, 33, 21], embodied AI [31, 106], and autonomous driving [84, 47, 77], where world models simulate environment dynamics and support downstream control and generalization. In such contexts, precise and interpretable evaluation is critical not only for benchmarking, but also for diagnosing failure modes and ensuring alignment with intended behaviors.
[…]
@@ Appendix B Reproducibility Statement
To support reproducibility and facilitate future research, we provide detailed instructions for reproducing all main experiments. Detailed descriptions of model architectures, training procedures, and dataset construction are provided in Section 4 and Appendix D. All experiments have been repeated for three runs. Plots and tables with quantitative results show the standard deviation across these runs.
[…]
@@
Appendix C Dataset
[…]
C.1 Construction Process
@@
    Description Generation. Next, for each sequence of frames $f^{(1:L)}\in\mathcal{V}$, we use the associated control log $c^{(1:L)}$ to extract action information and the metadata $m$ to obtain character-related attributes. These are combined to generate a structured natural language description via $d=\texttt{Describe}(c^{(1:L)},m)$. This yields a set of paired video–text examples: $\mathcal{Z}=\left\{(f^{(1:L)},d)\mid f^{(1:L)}\in\mathcal{V}\right\}$.
@@
    Question-Answer Pair Construction. Finally, we generate six QA pairs per clip, spanning two predefined tasks (AR and CR), each instantiated in three question formats: binary, multiple-choice, and open-ended. To enable this, we define task-specific answer spaces using $\texttt{GetAnswerSpace}(\mathcal{Z})$, which returns $\mathcal{Y}_{\text{AR}}$ for action categories and $\mathcal{Y}_{\text{CR}}$ for character identities, based on all video–text pairs in $\mathcal{Z}$. For each clip, we extract the task-specific ground-truth answer from the corresponding description as $y=\texttt{ExtractLabel}(d,t)$, where $t\in\{\text{AR},\text{CR}\}$. Each QA format is constructed as follows: (i) Binary: Two binary question-answer pairs are generated per instance using FormatBinaryPrompt. The positive question $Q^{\text{pos}}$ is constructed using the correct label $y\in\mathcal{Y}^{(t)}$ and paired with the positive answer $A^{\text{pos}}$. The negative question $Q^{\text{neg}}$ is constructed using an incorrect label $\tilde{y}\sim\texttt{SampleDistractor}(\mathcal{Y}^{(t)}\setminus\{y\})$ and paired with the negative answer $A^{\text{neg}}$. (ii) Multiple-Choice: A question $Q$ is generated using the full set of candidate options, formatted via FormatOptions$(\mathcal{Y}_{t})$. The question is constructed with FormatMCPrompt$(t,O)$ and paired with the correct answer $y\in\mathcal{Y}_{t}$. (iii) Open-Ended: A free-form question $Q$ is generated using FormatOEPrompt$(t)$, prompting the model to produce the correct label $y\in\mathcal{Y}_{t}$ without access to predefined answer choices.
@@
Appendix D Experimental Details
@@
D.1 Model
This section provides extended details on the architecture, pretraining configuration, and input formatting of the vision-language models used in our experiments. Our primary backbone is PaliGemma [10, 91].
@@ D.1 Model
D.1.1 Overview
[…]
@@ D.1.1 Overview
Language Decoder: Gemma. The language module $\mathcal{M}_{L}$ is a decoder-only Transformer with 26 layers and 2304-dimensional hidden states. Token embeddings are learned over a vocabulary of 257,216 tokens, encoded using the SentencePiece tokenizer [58]. Each Transformer block contains a self-attention mechanism with separate linear projections for queries, keys, and values. The MLP block follows a gated architecture, where the input is processed through parallel down projection and gating projection layers, modulated by a GELU-Tanh activation [44], combined via elementwise multiplication, and then passed through an up projection to return to the model’s hidden dimension. RMSNorm is applied before and after both attention and MLP sublayers to stabilize training. Rotary positional embeddings are added to enable relative position encoding. Output tokens are produced via a tied language modeling head that projects back to the vocabulary space.
D.1.2 Configurations

### v2

@@ 4 Experiments
We evaluate universe on ground truth video data, focusing on AR and CR across binary, multiple-choice, and open-ended formats. Our goals are twofold: (i) to benchmark performance against zero-shot and fine-tuned baselines, and (ii) to assess the trade-offs between adaptation strategies under constrained supervision and compute.
Figure 2: Comparison of universe and baseline models on Action and Character Recognition, all models trained for 1 epoch. Left: universe outperforms all baselines on AR. Right: On CR, it ranks third, behind models with either full vision encoder tuning or task-specific training with greater supervision. Trained under a unified protocol with minimal parameter updates (0.07%) and reduced per-task data, universe delivers strong performance across both tasks.
@@
[…]
@@ 5 Analysis
We find a consistent performance gap between AR and CR, highlighting the greater temporal and causal complexity of action understanding. This motivates our focus on AR as the more challenging and diagnostic task. Below, we analyze how adaptation choices shape performance on AR.
@@
Results. CR converges rapidly, achieving over 97% exact match after 12.5% of an epoch (~4K samples; Figure 5, bottom), and shows minor improvement with further training. In contrast, AR improves only modestly under extended training when limited to a single frame (Figure 5, top), suggesting that supervision alone is insufficient in the absence of temporal information. Motivated by this, we jointly scale both supervision and input length, varying the number of frames and epochs. As shown in Figure 3, performance on AR improves consistently across all formats, with the best results achieved under combined scaling.
@@
Temporal Sampling Strategies. Following the observation that AR requires both extended supervision and temporally rich input, we examine how frame selection impacts performance. We compare first-$n$ sampling, which selects the first $n$ consecutive frames from each rollout, to uniform-$n$ sampling, which draws $n$ evenly spaced frames across the entire sequence. We conduct experiments at varying context lengths, using $n\in\{1,2,\dots,8\}$ frames. to evaluate the impact of both sampling method and input horizon.
Figure 5: Exact Match accuracy for Action Recognition and Character Recognition.
@@
Optimizing Data Mix for Unified Multi-Task Evaluation. We analyze how training data composition affects multi-task performance in universe, with the goal of enabling a single model to generalize across AR and CR. Specifically, we study how the task-level ratio $\alpha$ and format-level ratio $\beta$ influence performance across evaluation settings. We first conduct a hierarchical ablation to identify an optimized data mixture, then assess its impact by comparing against a default task mix with uniform sampling.
@@ 6 Evaluating World Model Rollouts with universe
We evaluate the reliability of universe through a human study spanning eight distinct settings that vary in model scale, training data diversity, and output resolution. Our analysis considers two axes: (i) in-domain accuracy, measured on Skygarden, and (ii) generalization to six previously unseen environments.
Concretely, we study rollouts generated by: (i) a large-scale model trained across multiple environments with higher-resolution rollouts ($300\times 180$), and (ii) a smaller model trained on a single environment with lower-resolution rollouts ($128\times 128$). These two model families expose complementary challenges: resolution mismatch, domain coverage, and rollout fidelity. We construct eight evaluation settings: settings 1–7 draw from the large-scale model across diverse environments, while setting 8 uses the smaller model in the fine-tuning environment. Each setting contains 30 rollouts, paired with six natural-language questions from our evaluation protocol, yielding 240 rollouts in total. universe answers each question via majority vote over five greedy decoding samples. Human annotators rate responses on a four-point ordinal scale (Correct, Partially Correct, Incorrect, Unclear), with double annotation and adjudication on disagreements. Inter-rater reliability is measured using Cohen’s $\kappa$. Full details of the annotation protocol are provided in Appendix F.
Results. Figure 9 reports graded accuracy across all settings. Rollouts from the smaller, single-environment model yield lower evaluation accuracy, likely due to resolution mismatch, while the larger, multi-environment model provides higher-quality inputs. These results demonstrate that universe generalizes across model scales, rollout fidelities, and environments, while remaining closely aligned with human judgments.
Figure 8: Example frames from the eight evaluation settings, spanning different model scales, rollout fidelities, and environments. Note the resolution difference: Settings 1-7 feature $300\times 180$ rollouts, whereas Setting 8 uses $128\times 128$.
Figure 9: Graded accuracy of universe across the eight evaluation settings. Performance improves with higher-fidelity rollouts and remains stable across unseen environments (2–7). Cohen’s $\kappa$ indicates substantial inter-rater agreement.
@@ 7 Conclusion
In this paper, we investigate the use of Vision-Language Models (VLMs) as automated evaluators for world model rollouts, addressing the fundamental challenge of fine-grained, temporally grounded evaluation. We introduce a structured evaluation protocol centered on action and character recognition tasks across binary, multiple-choice, and open-ended formats. To support this, we propose universe, a unified method for adapting VLMs to this setting through mixed supervision, efficient frame sampling, and lightweight fine-tuning. Our large-scale study demonstrates that universe matches the performance of all task-specific checkpoints using a single unified model and aligns closely with human judgments, establishing it as a scalable, semantics-aware evaluator for evaluating world models, particularly when explicit ground truth is unavailable or costly to obtain.
Limitations. While our experiments focus on simulated environments, chosen both for their ground-truth availability and their direct relevance to large video-game and interactive-entertainment industries Kanervisto et al. (2025), evaluating universe in real-world settings remains an important next step. Our protocol focuses on foundational semantic tasks, and extending it to cover higher-level reasoning represents an important direction for future work. Scaling universe to long-horizon rollouts is also challenging, as fine-grained reasoning becomes harder with larger visual context. Our frame-sampling analysis suggests that more intelligent sampling or hierarchical summarization could address this, and we plan to explore such strategies. Finally, as with all pretrained VLMs, universe may inherit dataset biases and exhibit reduced reliability on ambiguous cases. While we examined them partially during evaluation, a deeper investigation into bias propagation is an exciting research direction for future work.
Appendix
@@ Appendix A Broader Impact
As world models become integral to simulation, planning, and decision-making in interactive environments, evaluation remains a key bottleneck for both research progress and safe deployment. We address this challenge by introducing a unified, sample-efficient framework for evaluating world model rollouts using adapted VLMs, designed for fine-grained, temporally grounded, and semantically coherent assessment.
This capability has direct implications for high-impact domains such as neural game engines (Kanervisto et al., 2025; Guo et al., 2025; Gao et al., 2025; Chen et al., 2025), embodied AI (Du et al., ; Yang et al., 2024), and autonomous driving (Russell et al., 2025; Hu et al., 2023a; Ni et al., 2025), where world models simulate environment dynamics and support downstream control and generalization. In such contexts, precise and interpretable evaluation is critical not only for benchmarking, but also for diagnosing failure modes and ensuring alignment with intended behaviors.
By reducing dependence on human annotation and task-specific fine-tuning, universe offers a scalable alternative that lowers the computational and environmental costs of rollout evaluation. However, reliance on automated evaluators introduces risks: adapted VLMs may inherit biases from pretraining, struggle under distributional shift, or yield unreliable judgments in edge cases. These risks are amplified in safety-critical settings, where miscalibrated evaluations can propagate downstream errors.
@@ Appendix B Reproducibility Statement
To support reproducibility and facilitate future research, we provide detailed instructions for reproducing all main experiments. Detailed descriptions of model architectures, training procedures, and dataset construction are provided in Section 4 and Appendix E. A high-level overview of the overall implementation framework is included in Appendix C.1. All experiments have been repeated for three runs. Plots and tables with quantitative results show the standard deviation across these runs.
Use of Existing Assets. We experiment with a range of open-weight VLMs, including three PaliGemma variants (version 1 (3B) (Beyer et al., 2024) and version 2 (3B and 10B) (Steiner et al., 2024)), VideoLLaMA3 (2B, 7B) (Boqiang Zhang, 2025), and CLIP (Radford et al., 2021) with the following vision encoder configurations: ViT-B/32, ViT-B/16, ViT-L/14, and ViT-L/14 with $336\times 336$ resolution. universe is built on top of PaliGemma v2 (3B), using publicly released checkpoints for initialization. Further architectural and implementation details are provided in Appendix E.1. For our software stack, we use Matplotlib (Hunter, 2007) for plotting, NumPy (Harris et al., 2020) for data handling, openCV (Bradski, 2000), FFmpeg (Tomar, 2006) and PIL (Umesh, 2012) for video and image processing, and NLTK (Bird & Loper, 2004) for text processing. Parameter-efficient fine-tuning is implemented using the PEFT library (Mangrulkar et al., 2022). We log our experiments using Weights and Biases (Biewald, 2020).
Use of Large Language Models. Portions of the manuscript were polished with the assistance of large language models (LLMs). The use of LLMs was limited to only improving readability and style; all ideas, and experimental designs were developed by the authors.
@@
Appendix C universe: Additional Details
C.1 Implementation Overview
This section outlines the implementation of universe in Python, presented as high-level pseudocode. The system is structured around two main stages: (i) Adaptation:fine-tuning a VLM on task-specific question-answer (QA) supervision derived from ground truth; (ii) Evaluation:using the adapted model to assess new rollouts via structured, prompt-based recognition tasks.
Adaptation Pipeline.
The adaptation stage can be implemented as two modules: AdaptationDatasetBuilder and VLMAdapter.
AdaptationDatasetBuilder.
This class constructs an adaptation dataset from raw ground truth data, initialized via load_ground_truth_data (see Section 3 and Appendix D). The core method, build, takes four arguments: alpha_task, which specifies the task mixture ratio; beta_format, which controls the distribution over QA prompt formats; context_length, which determines the number of frames per QA instance; and sampling_strategy, which defines how frames are sampled from rollouts. The builder first applies stratified_sample to select a subset of annotated samples that match the specified configuration. For each sample, it invokes _sample_visual_context to extract the relevant frames, and constructs a triplet consisting of frames, question, and answer.
VLMAdapter.
This class applies an adaptation strategy to a base VLM, passed via the base_vlm argument. Given an adaptation dataset adaptation_data, a tuning strategy specified by the strategy parameter, and a fixed number of training steps num_steps, the adapter trains the model by iteratively sampling a batch, computing the loss via compute_loss, and applying updates with update_model.
Evaluation Pipeline.
The evaluation stage can be implemented via two additional modules: RolloutsGenerator and Universe.
RolloutsGenerator.
This component autoregressively samples rollout trajectories from a world model (textttworld_model). Given an initial observation o_initial and an action sequence a_seq, the rollout method generates a sequence of predicted observations by maintaining lists of past observations (o_lt) and actions (a_lt). At each timestep, it calls predict_next_observation to obtain the next predicted frame, appends it to the rollout sequence o_seq, and continues until timestamps is reached. This process produces a full trajectory simulating environment dynamics.
Universe.
This module serves as the inference engine of our framework. It wraps an adapted VLM passed via adapted_vlm. Given a generated rollout and an evaluation specification, the method evaluate_rollout constructs a prompt using generate_question, parameterized by a recognition target and complexity level. It then calls evaluate, which queries the VLM with the resulting rollout and question, returning the model’s answer.
C.2 Adaptation to New Domains
To adapt universe to a new environment, one could collect a small set of reference trajectories that provide ground-truth observations and actions for the evaluation dimensions of interest. Question and answer templates can then be instantiated from these ground-truth signals to construct a mixed-supervision adaptation dataset.
In practice, two complementary routes exist for obtaining such reference data: (i) partnerships with game studios (as in the present work) provide high-quality logs and metadata but are resource-intensive, (ii) bootstrapping from publicly available gameplay video datasets, where generative models or synthetic data can supply initial labels.
At inference time, the evaluator could process rollouts by emitting natural-language responses along the specified dimensions. These outputs may be used either as structured feedback for improving the world model or mapped to numerical scores and aggregated into a single trajectory-level quality measure. Although our experiments use 14-frame inputs, this is a design choice rather than a conceptual constraint: the same procedure can extend to longer rollouts by increasing the visual context window or applying a sliding-window scheme over 14-frame segments.
Overall, deploying universe as a standalone evaluator would require: (i) a small adaptation dataset of reference trajectories with ground-truth labels, (ii) the task specification, and (iii) the lightweight fine-tuning recipe. Once adapted, the evaluator could provide dimension-wise assessments that can be aggregated or used as feedback in downstream world-model training.
Appendix D Dataset
This section details the construction and release of the dataset used to adapt VLMs for fine-grained evaluation of world model rollouts. We curate a realistic, human-centered dataset derived from actual gameplay in a complex multi-agent environment. Designed to provide temporally grounded and semantically structured supervision, the dataset aligns with the downstream evaluation setting and supports adaptation to both action and character recognition tasks across all QA formats. We describe the data construction pipeline, QA generation process, and release format below.
D.1 Construction Process
@@
    Description Generation. Next, for each sequence of frames $f^{(1:L)}\in\mathcal{V}$, we use the associated control log $c^{(1:L)}$ to extract action information and the metadata $m$ to obtain character-related attributes. These are combined to generate a structured natural language description via $d=\texttt{Describe}(c^{(1:L)},m)$, where Describe is a rule-based procedure that transforms the logged actions and metadata into textual descriptions used for constructing the QA supervision. This yields a set of paired video–text examples: $\mathcal{Z}=\left\{(f^{(1:L)},d)\mid f^{(1:L)}\in\mathcal{V}\right\}$.
@@
[…]
@@
Appendix E Experimental Details
@@
E.1 Model
This section provides extended details on the architecture, pretraining configuration, and input formatting of the vision-language models used in our experiments. Our primary backbone is PaliGemma (Beyer et al., 2024; Steiner et al., 2024).
@@ E.1 Model
E.1.1 Overview
PaliGemma is a VLM that processes both images and text as input and autoregressively generates natural language output. It follows the training paradigm of PaLI-3 (Chen et al., 2023), combining a ViT-based vision encoder (Dosovitskiy et al., 2021) with a decoder-only Transformer language model. The model is publicly available (Wolf et al., 2019). The architecture is fully modular, comprising three parameterized components: (i) Vision encoder($\mathcal{M}_{V}$): based on SigLIP (Zhai et al., 2023), specifically the “shape optimized” So400m (Alabdulmohsin et al., 2023). (ii) Multimodal projection head($\mathcal{M}_{P}$): a single linear layer for projecting visual features into the language decoder’s embedding space. (iii) Language decoder($\mathcal{M}_{L}$): a Transformer-based autoregressive model from the Gemma family (Mesnard et al., 2024; Rivière et al., 2024). Below, we discuss the architecture in more details, the general layer-level overview is also provided in Table 1.
@@ E.1.1 Overview
Language Decoder: Gemma. The language module $\mathcal{M}_{L}$ is a decoder-only Transformer with 26 layers and 2304-dimensional hidden states. Token embeddings are learned over a vocabulary of 257,216 tokens, encoded using the SentencePiece tokenizer (Kudo & Richardson, 2018). Each Transformer block contains a self-attention mechanism with separate linear projections for queries, keys, and values. The MLP block follows a gated architecture, where the input is processed through parallel down projection and gating projection layers, modulated by a GELU-Tanh activation (Hendrycks & Gimpel, 2016), combined via elementwise multiplication, and then passed through an up projection to return to the model’s hidden dimension. RMSNorm is applied before and after both attention and MLP sublayers to stabilize training. Rotary positional embeddings are added to enable relative position encoding. Output tokens are produced via a tied language modeling head that projects back to the vocabulary space.
E.1.2 Configurations